OpenStudy (anonymous):

write the standard form of the equation of the parabola whose directrix is x=-1 and whose focus is at (5,-2)

OpenStudy (anonymous):

\[ (x+1)^2 = (x-5)^2 + (x+2)^2 \]

OpenStudy (anonymous):

hmm not one of my choices

OpenStudy (anonymous):

\[ (x+1)^2 = (x-5)^2 + (y+2)^2 \]

OpenStudy (anonymous):

a) (y+2)^2=12(x+2) b) y-2=12(x+2)^2 c) x+2=1/12(y+2)^2 d) x-2=1/12(y+2)^3

OpenStudy (anonymous):

*****d)x-2=1/12(y+2)^2

OpenStudy (anonymous):

It is d. You can simplify my answer above to get that.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

\[ (x+1)^2-(x-5)^2=(y+2)^2\\ 12 x-24=(y+2)^2\\ 12( x-2)=(y+2)^2\\ ( x-2)=\frac 1 {12}(y+2)^2 \]